OpenStudy (qaisersatti):

please help me with Q5 gain and Q6

OpenStudy (anonymous):

Solution for Q5: Vout=Vs-(f*(Vgs^3))*R Gain = -3*f*(Vgs^2)*R

OpenStudy (qaisersatti):

thanks bro... can u help me with Q 3 if u have done it too..

OpenStudy (anonymous):

sorry, I did that one by simulation with Simulink :)

OpenStudy (qaisersatti):

and wat abt Q 6...do u have the formulas

OpenStudy (anonymous):

working on them, but I'm stuck on that, sorry!

OpenStudy (qaisersatti):

its ok i am also trying ...thanks brother